Let X be a random variable with density function
f(X) = {2/x^3, if x≥1
{0, otherwise
Show that E(X) exists and E(X) = 2 but Var(X) does not exist.
1
Expert's answer
2019-04-09T10:41:28-0400
"f(x)= \begin{matrix}
 2/x^3 & x\geqslant1, \\ 
 0, & otherwise
\end{matrix}"

"E(X)=\displaystyle\int_{1}^\infin x{2 \over x^3}dx=[{-2 \over x}]\begin{matrix}
 \infin \\
 1
\end{matrix}=2"

"Var(X)=E(X^2)-(E(X))^2"

"\displaystyle\int_{1}^\infin x^2{2 \over x^3}dx=2[ln|x|]\begin{matrix}
 \infin \\
 1
\end{matrix}=\infin"

Therefore, Var(X) does not exist.